From nch@roe.ac.uk Mon Dec 20 17:33:47 2004 Date: Fri, 10 Dec 2004 11:30:02 +0000 (GMT) From: Nigel Hambly To: WFCAM Science Archive Team -- Eckhard Sutorius , Harvey MacGillivray , Ian Bond , Mike Read , Nigel Hambly , Bob Mann Cc: CCs for WSA weekly meeting minutes distribution -- Clive Davenhall , Andrew Lawrence , Andy Adamson , Peter Shillan , John Taylor , Jim Emerson , Malcolm Stewart , Martin Hill , Mike Irwin , Peredur Williams Subject: WFAU WSA weekly meeting minutes: 10th December 2004 Minutes of WFCAM Science Archive meeting: 10th December 2004 ------------------------------------------------------------- ------------------------------------------------------------- Present: NCH, PMW, ETWS (lowest attendance ever!) Apologies: JPE, RGM, MCH, AL, HMG, JDT, GPS, MAR, JMS DONM: 10am, Friday 17th December, plate library Actions discharged: ------------------- ACTION: PMW to apply a little pressure to the premises team to solve the continuing problem of heat overload in C1. Discharged; new unit has been ordered; anticipate delivery and installation within the next week or so. ACTION: JPE to design and set up centralised VDFS web pages. Progressing. (there is now a link to it from our WSA Twiki). Discharged (should have been noted as discharged ages ago - sorry). ACTION: NCH, JMS & PMW to meet Wednesday 8th Dec at 10am in the Plate Library to thrash out the VDFS top level forward-plan of work. Discharged; new action (see below). ACTION: NCH to contact Eclipse to arrange reconfiguration of old catalogue server ahmose. Discharged; work is scheduled for the 20th December. Actions partly discharged but continuing: ----------------------------------------- ACTION: MAR to check out SQL Server networked server functionality for cross-server querying. Continues; NCH suggested we should do some experiments with the linked SQL Servers ahmose & amenhotep once the dust has settled on their reorganisation over the next couple of weeks. ACTION: NCH to remember to action a team programmer to implement range checking in the low-level ingest codes. ETWS & NCH have been discussing this; likely that the new developer will be assigned this task to start them off. ACTION: JDT to liaise with MCH concerning an end-to-end test of the PAL deployment. Continues; see Astrogrid Deployment below. - continuing; thanks for progressing these. Actions carried forward from 3/12/04 meeting: ---------------------------------------------- None this week. Specific points and new actions: -------------------------------- Project management: VDMT telecon meeting passed off without incident on Monday. Some discussions concerning the UK review (variously called PDR and FDR according to whim) and the associated timescale for production of documentation - as yet no firm dates have been set. NCH, PMW and JMS eventually met on Wednesday to dicsuss the top-level workpackage breakdown. ACTION: PMW to draft, and then iterate with NCH, the top-level WP breakdown for VDFS versions 2-4, consistent with VEGA and existing detailed work breakdown structure. WFCAM update: News from Hawaii indicates that WFCAM engineering has gone well, but there is still a lot to do before the camera goes back on the telescope in March '05, including larger than anticipated optical adjustments and some electronics work. the 2nd commissioning period in March is anticipated to take 1 to 2 weeks, but useful data may be obtained at that time. Comments and issues arising from CASU fortnightly minutes: No new minutes as of 10/12/04 am. Networking: Nothing new to report. Hardware: Very good news this week: the newly deployed catalogue server is behaving perfectly. One SCSI disk failure over last weekend did not affect archive availability at all; the offending disk was hot-swapped out on Tuesday and a new disk inserted, with RAID array automatically building and initialising, without any archive downtime whatsoever. Benchmark tests of the new server show that it is slightly slower for a single process, but that it is 3x faster relative to the single/muliple performance ratio of the old server, when executing 16 concurrent trawl queries (probably something to do with controller caching). Hot-swapping of RAID disks is extremely straight forward, so the large disk arrays will be easily maintainable by archive operator(s) when the hardware comes off contract maintenance, provided a small pool of spares is held. Software: ETWS reported continuing integration work on CASU tools (new versions have been received from CASU and checked into the WFAU CVS; MJI has also provided a URL link to CASU's own deployment of those tools in a database-driven, user tunable image combination service). ETWS has also investigated some EpyDoc issues with third-party middleware (no easy solution, but no big problem). NCH has been working steadily on CU19, testing out the preparation, backup and release procedure for prepared survey products. SSA: No news this week. Astrogrid deployment: NCH & MCH switched the PAL deployment over to the new server yesterday - seems to work fine (some issues over speed, but we suspect these are not server-side since the interactive SSA interface is fast). The AG deployment will now be left alone until after the AG consortium demo meeting next week. Miscellaneous: Nothing else this week.